OpenStudy (mathhelp346):

Please help: Write an equation in slope-intercept form for the line that contains (5,3) and is perpendicular to the line -x-4y=4x-y

OpenStudy (anonymous):

A parallel line has the same slope, a perpendicular line has a negative inverse slope.

OpenStudy (mathhelp346):

Ok

OpenStudy (anonymous):

Solve for y to change the equation to slope intercept form. That will tell you what the slope is.

OpenStudy (mathhelp346):

-5x/3=y?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Can you identify the slope?

OpenStudy (mathhelp346):

Is it -5x/3?

OpenStudy (anonymous):

No, the slope is a number without a variable.

OpenStudy (anonymous):

y=mx+b m=slope

OpenStudy (mathhelp346):

Ok can you help me find the slope

OpenStudy (anonymous):

\(y=-\frac{5}{3}x\)

OpenStudy (mathhelp346):

Yeah that's what I said

OpenStudy (anonymous):

The slope is the coefficient of the x term.

OpenStudy (mathhelp346):

Ohh ok

OpenStudy (anonymous):

The use the point-slope formula to write the new equation.

OpenStudy (mathhelp346):

So y-3=5/3(x-5)?

OpenStudy (anonymous):

Almost. The perpendicular line has a negative inverse slope.

OpenStudy (mathhelp346):

Ok so it's y-3=3/5(x-5)

OpenStudy (anonymous):

Yes. The question asks for the slope-intercept form, so solve for y.

OpenStudy (mathhelp346):

So the answer is y=3/5x?

OpenStudy (anonymous):

You got it!

OpenStudy (mathhelp346):

Ok thank you :)

OpenStudy (anonymous):

You're Welcome.